If x + y = k is normal to y2 = 12 x, then k is :
(A) 3                                        (B) 9
(C) –9                                      (D) –3

Grade:12

1 Answers

Ramesh V
14 years ago

slope of parabola at any point is :

dy/dx = 12/2y

slope of line is -1,

as line is normal, (-1)*(12/2y) = -1

y = 6 and so x is 3

so eqn of line is x+y=9

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult
to understand - post it here and we will get you the answer and detailed solution very quickly.We are all
IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Ramesh
IIT Kgp - 05 batch